Find the area of the region enclosed by x = y^2 - 4y + 2 and x = y-2
1
Expert's answer
2020-09-03T18:05:55-0400

"S=\u222b_1^4(y-2)dy -\u222b_1^4(y^2-4y+2)dy =\u222b_1^4(y-2) -(y^2-4y+2)dy=\u222b_1^4(y-2-y^2+4y-2)dy=\u222b_1^4(-y^2+5y-4)dy=(-y^3\/3+5y^2\/2-4y)|_1^4=-4^3\/3+5*4^2\/2-4*4-(-1^3\/3+5*1^2\/2-4*1)=-64\/3+40-16-(-1\/3+5\/2-4)=-64\/3+40-16+1\/3-5\/2+4=-21-3\/2+28=7-2.5=4.5"